# Characterizations of amenability for noncommutative dynamical systems and Fell bundles

### v1

@@ Abstract.
We prove that for a locally compact group $G$, a $\mathrm{C}^{*}$-dynamical system $(A,G,\alpha)$ is amenable in the sense of Anantharaman-Delaroche if and only if, for every other system $(B,G,\beta)$, the diagonal system $(A\otimes_{\max}B,G,\alpha\otimes^{d}_{\max}\beta)$ has the weak containment property (wcp).
@@ 1. Introduction
Amenability is a central notion connecting group theory, operator algebras and noncommutative dynamics. Since the foundational work of Anantharaman-Delaroche on amenable actions of locally compact groups on $\mathrm{C}^{*}$- and $\mathrm{W}^{*}$-algebras [7, 8], a rich theory has developed relating amenability to approximation properties, tensorial behaviour and structural properties of crossed products. The goal of this paper is to provide new, unified characterizations of amenability both for $\mathrm{C}^{*}$-dynamical systems and for Fell bundles over locally compact groups, and to exploit those characterizations to obtain permanence and nuclearity results.
@@
    $\alpha$ is amenable in the sense of Anantharaman-Delaroche (equivalently, the enveloping $\mathrm{W}^{*}$-action $\alpha^{\prime\prime}$ on $A^{\prime\prime}_{\alpha}$ is $\mathrm{W}^{*}$-amenable).
@@ 2.1. Preliminaries
As in [7], we say that a $\mathrm{W}^{*}$-action $\gamma$ of $G$ on $M$ is amenable (or $\mathrm{W}^{*}$-amenable) if there exists an equivariant conditional expectation $P\colon L^{\infty}(G)\bar{\otimes}M\to M$, where we identify each $m\in M$ with $1\otimes m\in L^{\infty}(G)\bar{\otimes}M$ and the action of $G$ on $L^{\infty}(G)\bar{\otimes}M$ is the diagonal action induced by conjugation by the left regular representation $\lambda\colon G\to\mathbb{B}(L^{2}(G))$ and $\gamma$. It was first noticed in [7, 8] that the amenability of $\gamma$ is related to that of the restriction of $\gamma$ to the center of $M$, $\gamma|_{Z(M)}$. Anantharaman-Delaroche proved they are equivalent under separability conditions and the general case is a consequence of [11, Theorem 3.17].
@@
In Appendix A, we address the problem of constructing the (von Neumann) tensor product of possibly non-normal, equivariant, contractive completely positive (ccp) maps between $\mathrm{W}^{*}$-algebras. Our approach combines techniques from [8, Lemme 2.1], [20, Theorem 3], and [23, Theorem 4]. While these methods are likely familiar to experts, we were unable to find a precise reference in the literature.
@@
Fix a ccp extension $\psi\colon\mathbb{B}(L^{2}(G)\otimes Y)\to\mathbb{B}(X)$ of $\psi_{0}$ and name $D$ the multiplicative domain of $\psi$ [10, Proposition 1.5.7 & Definition 1.5.8]. Recall that
@@ 2.2. SIN groups and amenability
The forward implication is already known [11, Theorem 7.2], but we provide a proof here for the convenience of the reader. Take any $\mathrm{C}^{*}$-algebra $B$ and equip it with the trivial action $\beta$ of $G$. Then $\alpha\otimes_{\max}^{d}\beta=\alpha\otimes^{d}\beta$ (in particular, $\alpha$) has the wcp, and
@@ 3. Amenability of Fell bundles and $\mathrm{C}^{*}$-partial actions
When we say that $\mathcal{B}=\{B_{t}\}_{t\in G}$ is a Fell bundle we mean that the disjoint union $\bigsqcup_{t\in G}B_{t}$ forms a $\mathrm{C}^{*}$-algebraic bundle over $G$ in the sense of [16, Ch. VIII, 16.2], with fibers $B_{t}$. Following [3], we denote by $C_{c}(\mathcal{B})$, $L^{1}(\mathcal{B})$, $C^{*}(\mathcal{B})$, $C^{*}_{\mathop{\rm r}}(\mathcal{B})$, and $\mathbb{k}(\mathcal{B})$ the *-algebra of continuous cross-sections with compact support, the $L^{1}$ cross-sectional algebra, the full cross-sectional $\mathrm{C}^{*}$-algebra, the reduced cross-sectional $\mathrm{C}^{*}$-algebra, and the $\mathrm{C}^{*}$-algebra of kernels of $\mathcal{B}$, respectively.
@@
the canonical surjective -homomorphism given by the reduced representation [15, Definition 2.7]. We say that $\mathcal{B}$ has the weak containment property (wcp) if $\lambda^{\mathcal{B}}$ is injective, that is, if the full and reduced cross-sectional $\mathrm{C}^{*}$-algebras of $\mathcal{B}$ coincide. Exel and Ng referred to this property as amenability of the Fell bundle, but in order to avoid confusion with other notions of amenability used in this article, we adopt the terminology wcp instead (see Definition 3.10 and Remark 3.11).
Whenever we say that a Fell bundle has an approximation property in the sense of Exel and Ng, we mean one of those introduced in [15, Definition 3.6]; by [15, Theorem 3.9], each of them implies the wcp. The $L^{1}$-approximation property, introduced by Abadie in [1, Definition 4.9(3)], also implies the wcp [1, Theorem 4.11], and is formally weaker than the Exel–Ng approximation properties [1, Proposition 4.10]. In what follows, we shall prove that all these approximation properties are in fact equivalent.
@@
We use [16, VIII 3.8] to regard the multiplier algebra $\mathcal{M}(B_{e})$ of the fiber over the unit $e\in G$ as the $\mathrm{C}^{*}$-algebra of multipliers of $\mathcal{B}$ of order $e$. As shown in [1], the natural actions $\mathcal{M}(B_{e})\times\mathcal{B}\to\mathcal{B}$ and $\mathcal{B}\times\mathcal{M}(B_{e})\to\mathcal{B}$ are continuous. Given $\xi,\eta\in C_{c}(G,\mathcal{M}(B_{e}))$, we define a continuous map
@@
A net $\{\Phi_{i}\}_{i\in I}\subset\mathrm{CLB}(\mathcal{B})$ converges to $\Phi\in\mathrm{CLB}(\mathcal{B})$ on compact slices [15, Definition 3.4] if
@@
By [16, II 14.6], $\Gamma^{\prime}$ is dense in $C_{c}(\mathcal{B})$ for the inductive limit topology, hence also dense for both $\|\cdot\|_{\infty}$ and $\|\cdot\|_{1}$. Let $\|\cdot\|$ denote either of these norms. For $f\in C_{c}(\mathcal{B})$, $n\in\mathbb{N}$, $z_{1},\ldots,z_{n}\in\mathbb{C}$, $g_{1},\ldots,g_{n}\in\Gamma$, and $i\in I$, one has
@@
Assume $\alpha$ is a $\mathrm{C}^{*}$-partial action of $G$ on $A$, let $\mathcal{B}$ denote its semidirect product bundle and let $\beta$ be the natural action of $G$ on $\mathbb{k}(\mathcal{B})$. Then $\beta$ is a Morita enveloping action of $\alpha$ [3, Theorem 6.1], and Morita enveloping actions are unique up to Morita equivalence [3, Proposition 6.3]. In particular, if $\alpha$ is already a $\mathrm{C}^{*}$-action, then $\beta$ is Morita equivalent to $\alpha$. Since $\mathrm{C}^{*}$-amenability is invariant under Morita equivalence of $\mathrm{C}^{*}$-actions [11, Proposition 3.20], Definition 3.10 is consistent with the usual notion of amenability for $\mathrm{C}^{*}$-actions. More generally, a $\mathrm{C}^{*}$-partial action $\alpha$ is $\mathrm{C}^{*}$-amenable if and only if any (equivalently, all) of its Morita enveloping actions are $\mathrm{C}^{*}$-amenable.
Each of the equivalent conditions in Theorem 3.2 can be reformulated in terms of $\alpha$. For instance, for $\mathrm{C}^{*}$-actions the positive approximation property of Exel and Ng coincides with the AP of [11, Definition 4.3]. (To get the equivalence, one may apply Lemma 3.4 with $\Gamma\subset C_{c}(\mathcal{B})$ consisting of sections of the form $t\mapsto f(t)a\delta_{t}$, where $f\in C_{c}(G)$ and $a\in A$.) On the other hand, some other approximation properties – such as the QAP of [11, Definition 4.1] or the weak versions of the AP and the QAP – do not correspond to the translated versions of the properties in Theorem 3.2.
@@
Let $\mathcal{A}$ and $\mathcal{B}$ be weakly equivalent Fell bundles over the same group $G$, and let $\alpha$ and $\beta$ denote the respective actions on the $\mathrm{C}^{*}$-algebras of kernels. By [5, Theorem 4.2], $\alpha$ is Morita equivalent to $\beta$. Combining this with [11, Proposition 3.20] and Theorem 3.2, we obtain the chain of equivalences
@@ 3.1. Amenability and nuclearity
Let $\beta$ be the natural action of $G$ on $\mathbb{k}(\mathcal{B})$ and $\mathcal{C}$ its semidirect product bundle. Then $\mathcal{B}$ is equivalent to $\mathcal{C}$ and $C^{*}(\mathcal{C})=\mathbb{k}(\mathcal{B})\rtimes_{\beta}G$. By [3, Corollary 5.2], $\mathbb{k}(\mathcal{B})$ is nuclear. Since $\beta$ is $\mathrm{C}^{*}$-amenable, [11, Theorem 7.2] implies that $C^{*}(\mathcal{C})$ is nuclear. Using [6, Theorem 4.5] we get that $C^{*}(\mathcal{B})$ is nuclear. ∎
@@
By [6, Proposition 4.13], $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is Morita equivalent to $C^{*}_{\mathop{\rm r}}(\mathcal{C})=\mathbb{k}(\mathcal{B})\rtimes_{{\mathop{\rm r}}\beta}G$. Hence, $\mathbb{k}(\mathcal{B})\rtimes_{{\mathop{\rm r}}\beta}G$ is nuclear and [18, Theorem 3.4] implies that $\mathbb{k}(\mathcal{B})$ is nuclear (unfortunately, in that article nuclear $\mathrm{C}^{*}$-algebras are called amenable). Corollary 5.2 of [3] implies $B_{e}$ is nuclear. Finally, if $G$ is inner amenable, then [19, Corollary 6.6] implies $\beta$ is $\mathrm{C}^{*}$-amenable. ∎
@@ 4. An approximation property of Bédos and Conti
The following definition extends [12, Definition A.2] to Fell bundles over locally compact groups.
@@
Let $G^{d}$ be the group $G$ with the discrete topology, and set $\mathcal{B}^{d}:=\{B_{t}\}_{t\in G^{d}}$. As sets, $\mathcal{B}$ and $\mathcal{B}^{d}$ coincide. By [12, Theorem A.4], there exist a *-representation $T^{d}\colon\mathcal{B}^{d}\to\mathbb{B}(Y)$ and an operator $V\in\mathbb{B}(X,Y)$ such that $Y=\overline{\mathop{\rm span}}\,T^{d}(\mathcal{B}^{d})VX$ and $\phi(b)=V^{*}T^{d}_{b}V$ for all $b\in\mathcal{B}^{d}$.
@@ 4.2. Ccp maps and amenability
Take a nondegenerate *-representation $T\colon\mathcal{B}\bigotimes_{\max}\mathcal{C}\to\mathbb{B}(X)$ whose restriction to any fiber is isometric. If $\tilde{T}\colon C^{*}(\mathcal{B}\bigotimes_{\max}\mathcal{C})\to\mathbb{B}(X)$ is the integrated form of $T$, then $\tilde{T}\circ\Psi$ is cp and it has a Stinespring dilation $\tilde{S}\colon C^{*}(\mathcal{A}\bigotimes_{\max}\mathcal{C})\to\mathbb{B}(Y)$ with associated map $V\in\mathbb{B}(X,Y)$. Naturally, we name $S\colon\mathcal{A}\bigotimes_{\max}\mathcal{C}\to\mathbb{B}(Y)$ the disintegrated form of $\tilde{S}$. Using [16, VIII 11.8] and the fact that $\Psi(f\oslash g)=(\Phi\circ f)\oslash g$ get get $V^{*}S_{f(s)\otimes g(t)}V=T_{\Phi(f(s))\otimes g(t)}$, for all $f\in C_{c}(\mathcal{A})$, $g\in C_{c}(\mathcal{C})$ and $(s,t)\in G\times H$. This implies $V^{*}S(A_{s}\otimes_{\max}C_{t})V\subseteq T(B_{s}\otimes_{\max}C_{t})$, then there exists a function $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}\colon\mathcal{A}\bigotimes_{\max}\mathcal{C}\to\mathcal{B}\bigotimes_{\max}\mathcal{C}$ such that $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(A_{s}\otimes_{\max}C_{t})\subseteq B_{s}\otimes_{\max}C_{t}$ and $T\circ\Phi(z)=V^{*}S_{z}V,$ for all $(s,t)\in G\times H$ and $z\in\mathcal{A}\bigotimes_{\max}\mathcal{C}$. This implies $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(a\otimes c)=\Phi(a)\otimes c$ and $\|\Phi(z)\|\leq\|V\|^{2}\|z\|=\|\tilde{\Phi}\|\|z\|=\|\Phi|_{A_{e}}\|\|z\|$. We have $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(f\oslash g)=(\Phi\circ f)\oslash g$, so [16, II 13.16] implies $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}$ is continuous and it is cp because $T\circ\Phi(z)=V^{*}S_{z}V$ (see Remark 4.8). Finally, $\|(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})|_{A_{e}\otimes_{\max}C_{e}}\|=\|\Phi|_{A_{e}}\otimes_{\max}\operatorname{id}_{C_{e}}\|=\|\Phi|_{A_{e}}\|$. ∎
@@
Let $\{\Phi_{i}\colon\mathcal{B}\to\mathcal{B}\}_{i\in I}$ be a net of ccp maps witnessing the BCAP of $\mathcal{B}$. Notice that $\Phi_{i}(\mathcal{B}_{H})\subseteq\mathcal{B}_{H}$ and that for all $n\in\mathbb{N}$ and $t\in H^{n},$ $\mathbb{M}_{t}(\mathcal{B})=\mathbb{M}_{t}(\mathcal{B}_{H})$. By [16, II 14.8], $\{f|_{H}\colon f\in C_{c}(\mathcal{B})\}=C_{c}(\mathcal{B}_{H})$. Then $\{\mathcal{B}_{H}\to\mathcal{B}_{H},b\mapsto\Phi_{i}(b)\}_{i\in I}$ is a net witnessing the BCAP of $\mathcal{B}_{H}$. ∎

### v2

@@ Abstract.
We prove that for a locally compact group $G$, a $\mathrm{C}^{*}$-dynamical system $(A,G,\alpha)$ is amenable if and only if, for every other system $(B,G,\beta)$, the diagonal system $(A\otimes_{\max}B,G,\alpha\otimes^{d}_{\max}\beta)$ has the weak containment property (wcp).
@@ 1. Introduction
Amenability is a central notion connecting group theory, operator algebras and noncommutative dynamics. In the setting of actions of locally compact groups on von Neumann algebras, amenability originates in Zimmer’s work on measurable dynamics and was studied systematically by Anantharaman-Delaroche; see [7, 8]. In this context, amenability is often referred to as Zimmer amenability.
For actions on $\mathrm{C}^{*}$-algebras, Anantharaman-Delaroche introduced an amenability notion for discrete groups via the bidual [9]. However, for a general locally compact group $G$ this approach does not extend directly, since the induced action on the bidual $A^{**}$ of a $G$-$\mathrm{C}^{*}$algebra AA need not be $w^{*}$-continuous. This conceptual difficulty was resolved more recently by Buss-Echterhoff-Willett [12] and, independently, by Ozawa-Suzuki [23], building on Ikunishi’s theory [19] of the enveloping $G$-$\mathrm{W}^{*}$-algebra $A^{\prime\prime}_{\alpha}$. These works show that several natural formulations of amenability for $\mathrm{C}^{*}$-dynamical systems of locally compact groups – including definitions via the enveloping $\mathrm{W}^{*}$-algebra and approximation properties – are equivalent.
Since these foundational developments, a rich theory has emerged relating amenability to approximation properties, tensorial behaviour, and structural features of crossed products. The goal of this paper is to provide new, unified characterizations of amenability both for $\mathrm{C}^{*}$-dynamical systems and for Fell bundles over locally compact groups, and to exploit those characterizations to obtain permanence and nuclearity results. Throughout the paper, the term amenable for $\mathrm{C}^{*}$-actions is used in this modern sense, and we freely rely on the equivalent characterizations established in the works cited above.
@@
    $\alpha$ is amenable (i.e., the enveloping $\mathrm{W}^{*}$-action $\alpha^{\prime\prime}$ on $A^{\prime\prime}_{\alpha}$ is $\mathrm{W}^{*}$-amenable in the sense of Anantharaman-Delaroche).
@@ 2.1. Preliminaries
As in [7], we say that a $\mathrm{W}^{*}$-action $\gamma$ of $G$ on $M$ is amenable (or $\mathrm{W}^{*}$-amenable) if there exists an equivariant conditional expectation $P\colon L^{\infty}(G)\bar{\otimes}M\to M$, where we identify each $m\in M$ with $1\otimes m\in L^{\infty}(G)\bar{\otimes}M$ and the action of $G$ on $L^{\infty}(G)\bar{\otimes}M$ is the diagonal action induced by conjugation by the left regular representation $\lambda\colon G\to\mathbb{B}(L^{2}(G))$ and $\gamma$. It was first noticed in [7, 8] that the amenability of $\gamma$ is related to that of the restriction of $\gamma$ to the center of $M$, $\gamma|_{Z(M)}$. Anantharaman-Delaroche proved they are equivalent under separability conditions and the general case is a consequence of [12, Theorem 3.17].
@@
In Appendix A, we address the problem of constructing the (von Neumann) tensor product of possibly non-normal, equivariant, contractive completely positive (ccp) maps between $\mathrm{W}^{*}$-algebras. Our approach combines techniques from [8, Lemme 2.1], [22, Theorem 3], and [25, Theorem 4]. While these methods are likely familiar to experts, we were unable to find a precise reference in the literature.
@@
Fix a ccp extension $\psi\colon\mathbb{B}(L^{2}(G)\otimes Y)\to\mathbb{B}(X)$ of $\psi_{0}$ and name $D$ the multiplicative domain of $\psi$ [11, Proposition 1.5.7 & Definition 1.5.8]. Recall that
@@ 2.2. SIN groups and amenability
The forward implication is already known [12, Theorem 7.2], but we provide a proof here for the convenience of the reader. Take any $\mathrm{C}^{*}$-algebra $B$ and equip it with the trivial action $\beta$ of $G$. Then $\alpha\otimes_{\max}^{d}\beta=\alpha\otimes^{d}\beta$ (in particular, $\alpha$) has the wcp, and
@@ 3. Amenability of Fell bundles and $\mathrm{C}^{*}$-partial actions
When we say that $\mathcal{B}=\{B_{t}\}_{t\in G}$ is a Fell bundle we mean that the disjoint union $\bigsqcup_{t\in G}B_{t}$ forms a $\mathrm{C}^{*}$-algebraic bundle over $G$ in the sense of [17, Ch. VIII, 16.2], with fibers $B_{t}$. Following [3], we denote by $C_{c}(\mathcal{B})$, $L^{1}(\mathcal{B})$, $C^{*}(\mathcal{B})$, $C^{*}_{\mathop{\rm r}}(\mathcal{B})$, and $\mathbb{k}(\mathcal{B})$ the *-algebra of continuous cross-sections with compact support, the $L^{1}$ cross-sectional algebra, the full cross-sectional $\mathrm{C}^{*}$-algebra, the reduced cross-sectional $\mathrm{C}^{*}$-algebra, and the $\mathrm{C}^{*}$-algebra of kernels of $\mathcal{B}$, respectively.
@@
the canonical surjective -homomorphism given by the reduced representation [16, Definition 2.7]. We say that $\mathcal{B}$ has the weak containment property (wcp) if $\lambda^{\mathcal{B}}$ is injective, that is, if the full and reduced cross-sectional $\mathrm{C}^{*}$-algebras of $\mathcal{B}$ coincide. Exel and Ng referred to this property as amenability of the Fell bundle, but in order to avoid confusion with other notions of amenability used in this article, we adopt the terminology wcp instead (see Definition 3.10 and Remark 3.11).
Whenever we say that a Fell bundle has an approximation property in the sense of Exel and Ng, we mean one of those introduced in [16, Definition 3.6]; by [16, Theorem 3.9], each of them implies the wcp. The $L^{1}$-approximation property, introduced by Abadie in [1, Definition 4.9(3)], also implies the wcp [1, Theorem 4.11], and is formally weaker than the Exel–Ng approximation properties [1, Proposition 4.10]. In what follows, we shall prove that all these approximation properties are in fact equivalent.
@@
We use [17, VIII 3.8] to regard the multiplier algebra $\mathcal{M}(B_{e})$ of the fiber over the unit $e\in G$ as the $\mathrm{C}^{*}$-algebra of multipliers of $\mathcal{B}$ of order $e$. As shown in [1], the natural actions $\mathcal{M}(B_{e})\times\mathcal{B}\to\mathcal{B}$ and $\mathcal{B}\times\mathcal{M}(B_{e})\to\mathcal{B}$ are continuous. Given $\xi,\eta\in C_{c}(G,\mathcal{M}(B_{e}))$, we define a continuous map
@@
A net $\{\Phi_{i}\}_{i\in I}\subset\mathrm{CLB}(\mathcal{B})$ converges to $\Phi\in\mathrm{CLB}(\mathcal{B})$ on compact slices [16, Definition 3.4] if
@@
By [17, II 14.6], $\Gamma^{\prime}$ is dense in $C_{c}(\mathcal{B})$ for the inductive limit topology, hence also dense for both $\|\cdot\|_{\infty}$ and $\|\cdot\|_{1}$. Let $\|\cdot\|$ denote either of these norms. For $f\in C_{c}(\mathcal{B})$, $n\in\mathbb{N}$, $z_{1},\ldots,z_{n}\in\mathbb{C}$, $g_{1},\ldots,g_{n}\in\Gamma$, and $i\in I$, one has
@@
Assume $\alpha$ is a $\mathrm{C}^{*}$-partial action of $G$ on $A$, let $\mathcal{B}$ denote its semidirect product bundle and let $\beta$ be the natural action of $G$ on $\mathbb{k}(\mathcal{B})$. Then $\beta$ is a Morita enveloping action of $\alpha$ [3, Theorem 6.1], and Morita enveloping actions are unique up to Morita equivalence [3, Proposition 6.3]. In particular, if $\alpha$ is already a $\mathrm{C}^{*}$-action, then $\beta$ is Morita equivalent to $\alpha$. Since $\mathrm{C}^{*}$-amenability is invariant under Morita equivalence of $\mathrm{C}^{*}$-actions [12, Proposition 3.20], Definition 3.10 is consistent with the usual notion of amenability for $\mathrm{C}^{*}$-actions. More generally, a $\mathrm{C}^{*}$-partial action $\alpha$ is $\mathrm{C}^{*}$-amenable if and only if any (equivalently, all) of its Morita enveloping actions are $\mathrm{C}^{*}$-amenable.
Each of the equivalent conditions in Theorem 3.2 can be reformulated in terms of $\alpha$. For instance, for $\mathrm{C}^{*}$-actions the positive approximation property of Exel and Ng coincides with the AP of [12, Definition 4.3]. (To get the equivalence, one may apply Lemma 3.4 with $\Gamma\subset C_{c}(\mathcal{B})$ consisting of sections of the form $t\mapsto f(t)a\delta_{t}$, where $f\in C_{c}(G)$ and $a\in A$.) On the other hand, some other approximation properties – such as the QAP of [12, Definition 4.1] or the weak versions of the AP and the QAP – do not correspond to the translated versions of the properties in Theorem 3.2.
@@
Let $\mathcal{A}$ and $\mathcal{B}$ be weakly equivalent Fell bundles over the same group $G$, and let $\alpha$ and $\beta$ denote the respective actions on the $\mathrm{C}^{*}$-algebras of kernels. By [5, Theorem 4.2], $\alpha$ is Morita equivalent to $\beta$. Combining this with [12, Proposition 3.20] and Theorem 3.2, we obtain the chain of equivalences
@@ 3.1. Amenability and nuclearity
Let $\beta$ be the natural action of $G$ on $\mathbb{k}(\mathcal{B})$ and $\mathcal{C}$ its semidirect product bundle. Then $\mathcal{B}$ is equivalent to $\mathcal{C}$ and $C^{*}(\mathcal{C})=\mathbb{k}(\mathcal{B})\rtimes_{\beta}G$. By [3, Corollary 5.2], $\mathbb{k}(\mathcal{B})$ is nuclear. Since $\beta$ is $\mathrm{C}^{*}$-amenable, [12, Theorem 7.2] implies that $C^{*}(\mathcal{C})$ is nuclear. Using [6, Theorem 4.5] we get that $C^{*}(\mathcal{B})$ is nuclear. ∎
@@
By [6, Proposition 4.13], $C^{*}_{\mathop{\rm r}}(\mathcal{B})$ is Morita equivalent to $C^{*}_{\mathop{\rm r}}(\mathcal{C})=\mathbb{k}(\mathcal{B})\rtimes_{{\mathop{\rm r}}\beta}G$. Hence, $\mathbb{k}(\mathcal{B})\rtimes_{{\mathop{\rm r}}\beta}G$ is nuclear and [20, Theorem 3.4] implies that $\mathbb{k}(\mathcal{B})$ is nuclear (unfortunately, in that article nuclear $\mathrm{C}^{*}$-algebras are called amenable). Corollary 5.2 of [3] implies $B_{e}$ is nuclear. Finally, if $G$ is inner amenable, then [21, Corollary 6.6] implies $\beta$ is $\mathrm{C}^{*}$-amenable. ∎
@@ 4. An approximation property of Bédos and Conti
The following definition extends [13, Definition A.2] to Fell bundles over locally compact groups.
@@
Let $G^{d}$ be the group $G$ with the discrete topology, and set $\mathcal{B}^{d}:=\{B_{t}\}_{t\in G^{d}}$. As sets, $\mathcal{B}$ and $\mathcal{B}^{d}$ coincide. By [13, Theorem A.4], there exist a *-representation $T^{d}\colon\mathcal{B}^{d}\to\mathbb{B}(Y)$ and an operator $V\in\mathbb{B}(X,Y)$ such that $Y=\overline{\mathop{\rm span}}\,T^{d}(\mathcal{B}^{d})VX$ and $\phi(b)=V^{*}T^{d}_{b}V$ for all $b\in\mathcal{B}^{d}$.
@@ 4.2. Ccp maps and amenability
Take a nondegenerate *-representation $T\colon\mathcal{B}\bigotimes_{\max}\mathcal{C}\to\mathbb{B}(X)$ whose restriction to any fiber is isometric. If $\tilde{T}\colon C^{*}(\mathcal{B}\bigotimes_{\max}\mathcal{C})\to\mathbb{B}(X)$ is the integrated form of $T$, then $\tilde{T}\circ\Psi$ is cp and it has a Stinespring dilation $\tilde{S}\colon C^{*}(\mathcal{A}\bigotimes_{\max}\mathcal{C})\to\mathbb{B}(Y)$ with associated map $V\in\mathbb{B}(X,Y)$. Naturally, we name $S\colon\mathcal{A}\bigotimes_{\max}\mathcal{C}\to\mathbb{B}(Y)$ the disintegrated form of $\tilde{S}$. Using [17, VIII 11.8] and the fact that $\Psi(f\oslash g)=(\Phi\circ f)\oslash g$ get get $V^{*}S_{f(s)\otimes g(t)}V=T_{\Phi(f(s))\otimes g(t)}$, for all $f\in C_{c}(\mathcal{A})$, $g\in C_{c}(\mathcal{C})$ and $(s,t)\in G\times H$. This implies $V^{*}S(A_{s}\otimes_{\max}C_{t})V\subseteq T(B_{s}\otimes_{\max}C_{t})$, then there exists a function $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}\colon\mathcal{A}\bigotimes_{\max}\mathcal{C}\to\mathcal{B}\bigotimes_{\max}\mathcal{C}$ such that $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(A_{s}\otimes_{\max}C_{t})\subseteq B_{s}\otimes_{\max}C_{t}$ and $T\circ\Phi(z)=V^{*}S_{z}V,$ for all $(s,t)\in G\times H$ and $z\in\mathcal{A}\bigotimes_{\max}\mathcal{C}$. This implies $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(a\otimes c)=\Phi(a)\otimes c$ and $\|\Phi(z)\|\leq\|V\|^{2}\|z\|=\|\tilde{\Phi}\|\|z\|=\|\Phi|_{A_{e}}\|\|z\|$. We have $(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})(f\oslash g)=(\Phi\circ f)\oslash g$, so [17, II 13.16] implies $\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}}$ is continuous and it is cp because $T\circ\Phi(z)=V^{*}S_{z}V$ (see Remark 4.8). Finally, $\|(\Phi\bigotimes_{\max}\operatorname{id}_{\mathcal{C}})|_{A_{e}\otimes_{\max}C_{e}}\|=\|\Phi|_{A_{e}}\otimes_{\max}\operatorname{id}_{C_{e}}\|=\|\Phi|_{A_{e}}\|$. ∎
@@
Let $\{\Phi_{i}\colon\mathcal{B}\to\mathcal{B}\}_{i\in I}$ be a net of ccp maps witnessing the BCAP of $\mathcal{B}$. Notice that $\Phi_{i}(\mathcal{B}_{H})\subseteq\mathcal{B}_{H}$ and that for all $n\in\mathbb{N}$ and $t\in H^{n},$ $\mathbb{M}_{t}(\mathcal{B})=\mathbb{M}_{t}(\mathcal{B}_{H})$. By [17, II 14.8], $\{f|_{H}\colon f\in C_{c}(\mathcal{B})\}=C_{c}(\mathcal{B}_{H})$. Then $\{\mathcal{B}_{H}\to\mathcal{B}_{H},b\mapsto\Phi_{i}(b)\}_{i\in I}$ is a net witnessing the BCAP of $\mathcal{B}_{H}$. ∎